# ImpScore: A Learnable Metric For Quantifying The Implicitness Level of Language

### v2

@@ Abstract
Handling implicit language is essential for natural language processing systems to achieve precise text understanding and facilitate natural interactions with users. Despite its importance, the absence of a metric for accurately measuring the implicitness of language significantly constrains the depth of analysis possible in evaluating models’ comprehension capabilities. This paper addresses this gap by developing a scalar metric that quantifies the implicitness level of language without relying on external references. Drawing on principles from traditional linguistics, we define “implicitness” as the divergence between semantic meaning and pragmatic interpretation. To operationalize this definition, we introduce ImpScore, a novel, reference-free metric formulated through an interpretable regression model. This model is trained using pairwise contrastive learning on a specially curated dataset comprising $112,580$ (implicit sentence, explicit sentence) pairs. We validate ImpScore through a user study that compares its assessments with human evaluations on out-of-distribution data, demonstrating its accuracy and strong correlation with human judgments. Additionally, we apply ImpScore to hate speech detection datasets, illustrating its utility and highlighting significant limitations in current large language models’ ability to understand highly implicit content. The metric model and its training data are available at https://github.com/audreycs/ImpScore.
@@ 2 Related Work
Linguistic Measures of Implicitness Theoretical linguistics has long examined how exactly meaning is conveyed through phenomena such as presupposition (Beaver, 1997) and implicature (Grice, 1975), without providing a quantitative measure of the implicitness of sentences. One of the few attempts at this goal, Vallauri & Masia (2014), proposed a system of quantification indexes to measure implicitness, focusing on concealment of content and concealment of responsibility as two core dimensions of implicitness. Garassino et al. (2022), on the other hand, proposed a taxonomy for classifying the implicit communication strategies of tweets. However, both studies rely heavily on manual identification of linguistic structures. Expert annotations would thus be required for scaling up their frameworks to large-scale datasets that could be used by modern NLP systems. This makes these two frameworks unsuitable for efficient and automatic evaluations of implicitness.
@@ 3.4 Integration and Model Embodiment
For model implementation, we embody the Text Encoder with Sentence-BERT (Reimers & Gurevych, 2019). with version all-mpnet-base-v2111https://sbert.net/docs/sentence_transformer/pretrained_models.html. We prioritize Sentence-BERT over other text encoders, such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), because Sentence-BERT is specifically finetuned to derive semantically meaningful sentence embeddings for sentence pairs, which we believe is beneficial for our task. The maximum input sequence length of Sentence-BERT is $384$ and its output dimension is $768$. We do not normalize the output embeddings. For the weights of matrices $\mathbf{W}_{p}$, $\mathbf{W}_{s}$, and $\mathbf{W}_{t}$, they are initialized with Xavier Initialization (Glorot & Bengio, 2010): $\mathbf{W}_{p},~{}\mathbf{W}_{s}\sim U[-\frac{\sqrt{6}}{\sqrt{d+l}},\frac{\sqrt{6}}{\sqrt{d+l}}],\mathbf{W}_{t}\sim U[-\frac{\sqrt{6}}{\sqrt{2l}},\frac{\sqrt{6}}{\sqrt{2l}}]$.
@@ 5.1 Experiment Setups
Hyperparameter Setting   We set the implicitness margin $\gamma_{1}$ to $0.5$, the pragmatic distance margin $\gamma_{2}$ to $0.7$, and $\alpha$ to $1.0$. Adam optimizer is used for gradient descent, with a batch size of $2^{13}$ and a learning rate of $0.01$. The data is randomly split into training, validation, and test sets in an $8:1:1$ ratio. We train for $30$ epochs, validating after each epoch, and save the model with the best validation performance on Implicitness Accuracy. The parameter size of ImpScore is about $105$ MB. We train ImpScore on one RTX A6000 GPU, costing around $1.5$ hours for one training.
@@ 5.2 Training Performance
Fig. 3 presents the performance of ImpScore on test set. The table in left panel shows the Implicitness Accuracy and Pragmatics Accuracy, along with the average implicitness scores for sentences ($I^{imp}$ and $I^{exp}$) and the average pragmatic distances for pairs ($\Delta P^{+}$ and $\Delta P^{-}$). ImpScore achieves over $95\%$ accuracy in both metrics. The center violin plot illustrates the distribution of implicitness scores of implicit sentences (blue), and for explicit sentences in positive pairs (green) and negative pairs (red). ImpScore effectively distinguishes between implicit and explicit ones, demonstrating a distributive bipolarity. The gaps between their average scores are even larger than $\gamma_{1}=0.5$. The right violin plot presents the distribution of pragmatic distances of positive pairs (green) and negative pairs (red). ImpScore effectively captures closer pragmatics in positive pairs. Although the output embeddings of ImpScore were not normalized, the results suggest that the range of pragmatic distances remains constrained within a narrow section. This is influenced by the properties of embeddings produced by Sentence-BERT.
@@ 6.1 OOD Data Preparation
We first prompted ChatGPT with $10$ different topics [^2] Please see the topic design, user study platform, and user participant recruitment details in Appendix A.5. shown in Tab. 4 and asked it to generate $4$ sentences with varying implicitness levels (most explicit, explicit, implicit, and most implicit) for each topic. The sentences in each topic formed a group. We, along with a linguist, then manually refined and verified the sentences in each group to ensure they manifest distinct levels of implicitness, allowing us to assign accurate gold ranks (ranked from most explicit to most implicit). The detailed sentences for each group are in Appendix A.3.
@@ 7.2 Language Models’ Capabilities on Different Implicitness Levels
8 Conclusion and Future Work
In this paper, we addressed the problem of quantifying the implicitness level of language. By providing an interpretable definition of “implicitness,” we have developed ImpScore, a metric capable of computing implicitness scores for sentences. ImpScore’s performance on out-of-distribution test data demonstrates satisfactory accuracy and alignment with human judgments, which highlights its generalizability. Furthermore, two applications underscored the effectiveness of ImpScore and revealed significant limitations in current LLMs in processing highly implicit content. These insights are crucial for advancing our understanding of the subtleties in language that LLMs currently struggle to grasp.
Despite these advancements, the research presented herein is not without its limitations, which open several avenues for future exploration. Firstly, while the dataset comprising $112,580$ sentence pairs has proven adequate, enlarging this dataset could enhance both the accuracy and the generalizability of ImpScore. Secondly, the current evaluation of ImpScore, though comprehensive, could be extended through additional user studies and a broader range of inventory questions. To this end, we encourage the research community to engage with ImpScore, exploring its applications and further refining its capabilities.
Looking forward, the potential applications of ImpScore in NLP research are vast and varied. It could be instrumental in measuring the implicitness level of machine-generated text, both in retrospective analyses and during the training phase of models. Specifically, ImpScore could serve as a valuable tool in reinforcement learning setups, providing a reward signal that enhances language models’ ability to handle nuanced linguistic constructs. Additionally, ImpScore’s capacity for screening complex, high-implicitness data could significantly benefit tasks in intent detection, social awareness, and theory of mind analyses, among others.
@@ A.5 User Study Design
OOD Data Topics We adopted most of the topics from the book Crucial conversations: tools for talking when stakes are high (Kohnen, 2008), section “Some Common Crucial Conversations” in Chapter One. This book introduces how to apply different expressions in important conversations, which we think is a very useful resource. Other topics were designed by ourselves. The details are shown in Tab. 11. We also asked a linguist to assess the quality of sentences under each group in Tab. 11. The quality of the test examples was evaluated on two dimensions: pragmatics and implicitness. The pragmatics rating (00 to $4$) reflects the level of pragmatic agreement among the sentences. If all four sentences convey the same meaning, the group receives a score of $4$. If three of the sentences share the same meaning, the group is given a score of $3$, and so on. In terms of the implicitness rating, since $\binom{4}{2}=6$ pairs can be created from the four sentences in a group, the implicitness rating will range from 00 to $6$, measuring the extent to which the four sentences show a progressively higher level of implicitness.

### v3

# ImpScore: A Learnable Metric For Quantifying The Implicitness Level of Sentence

@@ Abstract
Handling implicit language is essential for natural language processing systems to achieve precise text understanding and facilitate natural interactions with users. Despite its importance, the absence of a metric for accurately measuring the implicitness of language significantly constrains the depth of analysis possible in evaluating models’ comprehension capabilities. This paper addresses this gap by developing a scalar metric that quantifies the implicitness level of language without relying on external references. Drawing on principles from traditional linguistics, we define “implicitness” as the divergence between semantic meaning and pragmatic interpretation. To operationalize this definition, we introduce ImpScore, a reference-free metric formulated through an interpretable regression model. This model is trained using pairwise contrastive learning on a specially curated dataset consisting of (implicit sentence, explicit sentence) pairs. We validate ImpScore through a user study that compares its assessments with human evaluations on out-of-distribution data, demonstrating its accuracy and strong correlation with human judgments. Additionally, we apply ImpScore to hate speech detection datasets, illustrating its utility and highlighting significant limitations in current large language models’ ability to understand highly implicit content. Our metric is publicly available at https://github.com/audreycs/ImpScore.
@@ 2 Related Work
Linguistic Measures of Implicitness Theoretical linguistics has long examined how exactly meaning is conveyed through phenomena such as presupposition (Beaver, 1997) and implicature (Grice, 1975), without providing a quantitative measure of the implicitness of sentences. Among the few relevant works, Vallauri & Masia (2014), proposed a system of quantification indexes to measure implicit causality bias, focusing on how specific verbs influence causal attributions by calculating bias scores for each verb. Garassino et al. (2022) qualitatively assessed participants’ abilities to detect and interpret different types of implicit content. However, neither study addressed calculating implicitness scores for entire sentences. Additionally, their results heavily rely on human annotations, making them difficult to reproduce automatically.
@@ 3.4 Integration and Model Embodiment
For model implementation, we embody the Text Encoder with Sentence-BERT (Reimers & Gurevych, 2019) with version all-mpnet-base-v2111https://sbert.net/docs/sentence_transformer/pretrained_models.html. We prioritize Sentence-BERT over other text encoders, such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), because Sentence-BERT is specifically finetuned to derive semantically meaningful sentence embeddings for sentence pairs, which we believe is beneficial for our task. The maximum input sequence length of Sentence-BERT is $384$ and its output dimension is $768$. We do not normalize the output embeddings. For the weights of matrices $\mathbf{W}_{p}$, $\mathbf{W}_{s}$, and $\mathbf{W}_{t}$, they are initialized with Xavier Initialization (Glorot & Bengio, 2010): $\mathbf{W}_{p},\leavevmode\nobreak\ \mathbf{W}_{s}\sim U[-\frac{\sqrt{6}}{\sqrt{d+l}},\frac{\sqrt{6}}{\sqrt{d+l}}],\mathbf{W}_{t}\sim U[-\frac{\sqrt{6}}{\sqrt{2l}},\frac{\sqrt{6}}{\sqrt{2l}}]$.
@@ 5.1 Experiment Setups
Hyperparameter Setting   We set the implicitness margin $\gamma_{1}$ to $0.5$, the pragmatic distance margin $\gamma_{2}$ to $0.7$, and $\alpha$ to $1.0$. The dimension of two spaces is $64$. Adam optimizer is used for gradient descent. The data is randomly split into training, validation, and test sets in an $8:1:1$ ratio. We train for $30$ epochs, validating after each epoch, and save the model with the best validation performance on Implicitness Accuracy. The parameter size of ImpScore is about $105$ MB.
@@ 5.2 Training Performance
Fig. 3 presents the performance of ImpScore on test set. The table in left panel shows the Implicitness Accuracy and Pragmatics Accuracy, along with the average implicitness scores for sentences ($I^{imp}$ and $I^{exp}$) and the average pragmatic distances for pairs ($\Delta P^{+}$ and $\Delta P^{-}$). ImpScore achieves over $95\%$ accuracy in both metrics. The center violin plot illustrates the distribution of implicitness scores of implicit sentences (blue), and for explicit sentences in positive pairs (green) and negative pairs (red). ImpScore effectively distinguishes between implicit and explicit ones, demonstrating a distributive bipolarity. The gaps between their average scores are even larger than $\gamma_{1}=0.5$. The right violin plot presents the distribution of pragmatic distances of positive pairs (green) and negative pairs (red). ImpScore effectively captures closer pragmatics in positive pairs. Although the output embeddings of ImpScore were not normalized, the results suggest that the range of pragmatic distances remains constrained within a narrow section. This is influenced by the properties of embeddings produced by Sentence-BERT. A breakdown of ImpScore’s testing performance on different data sources is in Appendix A.11.
@@ 6.1 OOD Data Preparation
We prompted GPT-4 with $20$ different topics [^2] Please see the topic design, user study platform, and user participant recruitment details in Appendix A.5. to generate $4$ sentences with varying implicitness levels for each topic (most explicit, explicit, implicit, and most implicit). The sentences in each topic formed a group (i.e., $G_{i}$). We, along with a linguist, then manually refined and verified the sentences in each group to ensure they manifest distinct levels of implicitness, allowing us to assign accurate gold ranks (ranked from most explicit to most implicit). Here, we show the user study procedure and results for $10$ topics in Tab. 4. Results on the remaining $10$ topics are in Appendix A.13. The detailed sentences for each group in Tab. 4 are in Appendix A.3.
@@ 7.2 Language Models’ Capabilities on Different Implicitness Levels
Additionally, we applied ImpScore to assess the generation capabilities of different LLMs, as detailed in Appendix A.15.
@@
8 Conclusion and Discussion
In this paper, we addressed the problem of quantifying the implicitness level of language. By providing an interpretable definition of “implicitness,” we developed a metric capable of computing implicitness scores for sentences. ImpScore’s performance on out-of-distribution test data demonstrates satisfactory accuracy and alignment with human judgments, which highlights its generalizability.
However, this research has limitations that suggest directions for future work. First, while the dataset comprising $112,580$ sentence pairs, expanding it could improve ImpScore ’s accuracy and generalizability. To do so, beyond expanding the search for existing datasets, leveraging LLMs to generate synthetic data while maintaining a coherent speech style is a promising approach. Moreover, incorporating a broader range of implicitness types and multilingual data would enhance ImpScore ’s versatility. Second, ImpScore ’s evaluation, though comprehensive, could benefit from further user studies and a wider range of inventory questions. Third, as the dataset grows, employing more advanced text encoders like SimCSE (Gao et al., 2021) may further improve performance. To this end, we encourage the research community to explore and refine ImpScore ’s capabilities.
Looking forward, ImpScore has broad potential in NLP. It could assess implicitness in machine-generated text, aiding both retrospective analysis and model training. Specifically, ImpScore could serve as a tool in reinforcement learning setups, providing a reward signal that enhances language models’ ability to handle nuanced linguistic constructs. Additionally, ImpScore’s capacity for screening complex, high-implicitness data could significantly benefit tasks in intent detection, social awareness, and theory of mind analyses, among others.
Acknowledgements
This research was supported in part by a grant from the US National Library of Medicine (R01LM013833).
@@ A.5 User Study Design
OOD Data Topics We adopted most of the topics from the book Crucial conversations: tools for talking when stakes are high (Kohnen, 2008), section “Some Common Crucial Conversations" in Chapter One. This book introduces how to apply different expressions in important conversations, which we think is a very useful resource. Other topics were designed by ourselves. The details are shown in Tab. 11. We also asked a linguist to assess the quality of sentences under each group in Tab. 11. The quality of the test examples was evaluated on two dimensions: pragmatics and implicitness. The pragmatics rating (00 to $4$) reflects the level of pragmatic agreement among the sentences. If all four sentences convey the same meaning, the group receives a score of $4$. If three of the sentences share the same meaning, the group is given a score of $3$, and so on. In terms of the implicitness rating, since $\binom{4}{2}=6$ pairs can be created from the four sentences in a group, the implicitness rating will range from 00 to $6$, measuring the extent to which the four sentences show a progressively higher level of implicitness.
@@ A.10 User Study Questionnaire
A.11 Training Performances Breakdown
Table 13 includes the performance breakdown of ImpScore in different data sources in the test set during training.
Table 13: Breakdown of ImpScore’s performance on different data sources in the test set during training.
Source of Test Data Source Topics # (implicit, explicit) Pair # Correct Prediction Accuracy
DGHSID Implicit Hate Speech Detection $302$ $245$ $0.811$
Latent Hatred Implicit Hate Speech Detection $1166$ $1156$ $0.991$
ToxiGEN Implicit Hate Speech Detection $270$ $224$ $0.830$
Adversarial Implicit Hate Speech Detection $2518$ $2460$ $0.977$
SBIC Implicit Hate Speech Detection $900$ $896$ $0.996$
Implicatures Natural Language Inference $89$ $68$ $0.764$
ImpPres Natural Language Inference $12$ $12$ $1.000$
LUDWIG Natural Language Inference $69$ $55$ $0.797$
ABSA Sentiment Analysis $136$ $112$ $0.824$
SemEval Irony / Sarcasm Detection $132$ $108$ $0.818$
Snarks Irony / Sarcasm Detection $16$ $10$ $0.625$
R-3 Irony / Sarcasm Detection $14$ $12$ $0.857$
PDTB Irony / Sarcasm Detection $6$ $4$ $0.667$
A.12 Cross-lingual Performance of ImpScore
We conducted an experiment to evaluate ImpScore ’s performance on three non-English languages: French, German, and Chinese. Specifically, we first utilized Google Translate [^6] https://translate.google.com/ to translate the $10$ group topics from the user study in §6 into these non-English languages. Subsequently, we prompted ChatGPT in each non-English language to generate four sentences per topic, varying in levels of implicitness. Here are the prompt formats:
  1. 1.
    French: Pour ce sujet « Mettre fin à une relation », pourriez-vous générer 4 phrases avec différents niveaux d’implicit? du plus explicite au plus implicite.
  2. 2.
    German: Formulieren Sie zum Thema „Beziehung beenden“ bitte 4 Sätze mit unterschiedlichen Implizitheitsgraden, vom explizitesten bis zum implizitesten.
  3. 3.
    Chinese: {CJK}UTF8gbsn对于这个主题“结束关系。”,你能生成 4 个具有不同隐含程度的句子吗?(从隐含程度最低到最高)
The generated sentences in these languages were then translated back to English using Google Translate. We used Google Translate instead of directly using ChatGPT for translation to minimize inherent cross-language alignment biases in ChatGPT, ensuring more independence of each language. Then, we use ImpScore to rank the four sentences within each topic and compare the predictions with the gold rankings. The average results are summarized below. Detailed scores for each group are visible in the Tab. 14.
These findings indicate that ImpScore demonstrates a moderate positive correlation with gold rankings in these translated sentences. However, its accuracy diminishes compared to the assessment on original English sentences in our paper (avg. $\tau=0.76$ and $\rho=0.84$), particularly for language that is linguistically distant from English (i.e., Chinese).
This reduction in accuracy may stem from a loss of subtle implicitness nuances in original languages during translation into English, suggesting potential limitations in using translations as a proxy for cross-lingual evaluation of implicitness.
Table 14: ImpScore’s performance on translated text of non-English sentences.
Non-English Languages Avg. $\tau$ over $10$ groups Avg. $\rho$ over $10$ groups
French $0.63$ $0.74$
German $0.63$ $0.70$
Chinese $0.43$ $0.58$
A.13 Additional User Study Results
We conducted another $10$ groups OOD test data for user study, following the same procedure detailed in the paper in §6. Below are the topics:
  • $G_{1}$: Reject an auto dealerships offer
  • $G_{2}$: Talking to a coworker about a specific hygiene problem
  • $G_{3}$: A student cheating on the exam
  • $G_{4}$: Asking a friend to repay a loan
  • $G_{5}$: Complain about micromanaging
  • $G_{6}$: Unpleasant food smell
  • $G_{7}$: Criticizing a colleague’s work
  • $G_{8}$: Eating with big sound
  • $G_{9}$: Having bad memory about hometown
  • $G_{10}$: Talking to a team member who isn’t keep commitments
[…]
The results indicate a generally high accuracy of ImpScore with the gold ranks and concordant with human participants, which align with our conclusion in line 406-409 in the paper. In set 1, ImpScore slightly underperforms the average human ranking, and in set 2, ImpScore slightly surpasses the average human ranking.
Table 15: Performances of ImpScore and human participants compared to the gold ranks on additional $10$ data groups.
Set Human 1 Human 2 Human 3 Human 4 Human 5 ImpScore
1 $0.87$ $0.92$ $1.00$ $1.00$ $0.93$ $0.96$ $0.87$ $0.92$ $0.87$ $0.92$ $0.80$ $0.88$
2 $0.87$ $0.92$ $0.93$ $0.96$ $0.87$ $0.92$ $0.67$ $0.72$ $0.93$ $0.96$ $0.93$ $0.96$
Figure 19: $\tau$ correlation values between participants.
Figure 20: $\rho$ correlation values between participants.
A.14 Qualitative Analysis of ImpScore’s Performance
We evaluated the ImpScore’s performance on the in-distribution test set on the following features.
A.14.1 The length of the sentences
For $5,630$ pairs of (implicit sentence, explicit sentence) in the test data, we analyzed sentence length by calculating the average number of space-separated tokens per sentence. We then plotted the distribution of sentence lengths alongside the corresponding numbers of correctly predicted cases. The result plot is in Fig. 21, which shows no significant decline in accuracy for longer sentences.
Figure 21: ImpScore’s prediction performance on different length of sentence pairs.
A.14.2 Sentiment
We evaluated the sentiment of each sentence in the pairs using the pre-trained sentiment classification model cardiffnlp/twitter-roberta-base-sentiment-latest. Sentiments were categorized into {positive, neutral, negative}, and each pair was grouped into one of eight combinations (e.g., negative-positive, neutral-neutral).
The result plot is in Tab. 22. The “negative-positive" group exhibits the lowest performance, likely due to a bias introduced by its small sample size. Otherwise, the results do not reveal a clear underperformance for any specific sentiment group. Additionally, we observed that negative sentiment dominates the dataset—a trend that aligns with our expectations, as implicit sentences often arise in sensitive or negative contexts.
Figure 22: ImpScore’s prediction performance on different sentiment of sentence pairs.
A.14.3 TSNE for correct/incorrect predictions
We also analyzed the semantic distribution of correctly predicted pairs and incorrectly predicted pairs. To do so, we concatenated the implicit and explicit sentences in each pair, generated embeddings using Sentence-BERT, and applied t-SNE for dimensionality reduction to visualize the results in a 2D plot. The visualization is in Fig. 23.
The result reveals that while both correct and incorrect predictions are broadly distributed, certain clusters of incorrect predictions emerge (highlighted as Box 1 and Box 2 in the plot). On closer examination, we found that the incorrect predictions in Box 1 predominantly pertain to digital devices, such as “computers,” “CDs,” “laptops,” and “RAM.” This suggests a potential challenge for ImpScore in handling domain-specific or technical terminology. Box 2 contains pairs where distinguishing the implicitness levels between sentences proved exceptionally difficult. For example: (- Will that make everything alright? - Thank you., - Will that make everything alright? - Thank you for your concern.)
Figure 23: Embedding distributions of ImpScore’s correctly and incorrectly predicted sentence pairs.
A.15 Applying ImpScore to the generated sentences of LLMs
We zero-shot prompted seven LLMs (GPT-3.5, GPT-4o, Llama-3.2-Instruct, Mistral-7B, Mistral-Nemo, Claude-haiku, and Claude-sonnet) with $10$ topics same with the ones in the user study and asked them to generate four sentences with different implicitness levels. We used the prompt “Given the topic [topic], please generate four sentences about it, ranging from most explicit to most implicit. Output each sentence on a new line.” and set the temperature to each model’s default value.
For each run, we calculate the average implicitness scores of sentences over $10$ groups in each implicitness level (most explicit, explicit, implicit, most implicit). We conducted $5$ runs and plot the mean and standard deviation of the $5$ average scores. The result is in Tab. 24.
The results demonstrate that, according to ImpScore’s evaluation, all the models are generally capable of producing sentences with distinguishable implicitness levels. Among them, GPT-4o exhibited the best performance in generating sentences with the most clearly distinguishable levels of implicitness. Additionally, GPT-4o and Claude-sonnet showed the strongest ability to generate highly implicit sentences, aligning with our expectations given their advanced capabilities.
Figure 24: ImpScore’s scores on the generated sentences of seven LLMs.